OpenStudy (anonymous):

What is the simplified form of the expression 3[14divide(4^2-3^2)-5] A. -9 B. 27 C. 21 D. 6

OpenStudy (anonymous):

A

OpenStudy (anonymous):

You sure

OpenStudy (anonymous):

100%

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Would you like it expalined?

OpenStudy (anonymous):

Why not

OpenStudy (anonymous):

Plz explain so i know how to in the future

OpenStudy (anonymous):

\[3\left(\frac{ 14 }{4^2-3^2 }-5 \right)=3\left( \frac{ 14 }{ 7 } -5\right) =3\left( 2-5\right) =3(-3)\]

OpenStudy (anonymous):

Thank you:)